OpenStudy (anonymous):

Using the formula: a1/1-r (sum of an infinite geometric series) find the decimal: .7852852852852852... in fraction form.

OpenStudy (anonymous):

\[.7852852852852852 .... =\frac{785}{10^3}+\frac{285}{10^6}+\frac{285}{10^9}+\frac{285}{10^{1 2}}=\\ +\frac{785}{10^3}+\frac{285}{10^6}\left(1+\frac{1}{10^3}+\frac{1}{10^6}+\cdots\right) \]

OpenStudy (anonymous):

Use the following. If |r|<1, then\[ 1+r + r^2 + r^3 + \cdots = \frac 1{1-r} \] To sum the infinite geometric series in my previous post.

OpenStudy (anonymous):

You need to sum the infinite series \[ 1+\frac{1}{10^3}+\frac{1}{10^6}+\cdots \] Use the above formula in my previuous post with \[ r=\frac 1{10^3} \]

OpenStudy (anonymous):

\[.7852852852852852 .... =\frac{785}{10^3}+\frac{285}{10^6}+\frac{285}{10^9}+\frac{285}{10^{1 2}}+\cdots=\\ +\frac{785}{10^3}+\frac{285}{10^6}\left(1+\frac{1}{10^3}+\frac{1}{10^6}+\cdots\right) \]

OpenStudy (anonymous):

\[ 1+\frac{1}{10^3}+\frac{1}{10^6}+\cdots = \frac {1}{ 1-\frac 1{10^3}} =\frac{1000}{999} \]

OpenStudy (anonymous):

Finally \[\frac{785}{10^3}+\frac{285}{10^6} \frac{1000}{999} \] Simplify and you find your answer.

OpenStudy (anonymous):

It seems that you are not online\[ \frac{785}{10^3}+\frac{285}{10^6} \frac{1000}{999}=\frac{523}{666} \]

OpenStudy (anonymous):

In general any repeating decimals can be written as a fraction using the above method.